# Geometric Solution of the Loop Equation and Quark Confinement in Planar QCD

### v2

@@ 6 Self-Duality of the Area Derivative
$\ast\frac{\delta S_{\mathrm{phys}}}{\delta\sigma_{\mu\nu}}=\frac{\delta S_{\mathrm{phys}}}{\delta\sigma_{\mu\nu}}.$ (6.1)
@@
The physical action $S_{\mathrm{phys}}[\mathcal{C}]$ is the envelope (the maximum) of a family of functionals $S_{R}[C]:=S_{\chi}[RC;P]$, indexed by the rotation $R\in O(4)$, where $S_{\chi}[C;P]$ is the base functional from Sec. 4. The area derivative operator $\mathcal{D}:=\frac{\delta}{\delta\sigma}$ and the Hodge star $\ast$ are both $O(4)$-covariant. This means that if we can prove self-duality for any single member of this family (e.g., the $R=I$ case, $S_{\chi}[C;P]$), then self-duality holds for all members. The derivative of the envelope $S_{\mathrm{phys}}$ at a loop $C$ is, by definition, the derivative of the specific functional $S_{R_{\text{max}}}[C]$ that achieves the maximum at that point. Since all functionals in the family will be shown to have an SD derivative, $S_{\mathrm{phys}}$ must also have an SD derivative. Therefore, the proof reduces to verifying the self-duality for the base functional $S_{\chi}[C;P]$ (for $R=I$), which we do now.
@@
where $T_{\mu\nu}$ is the normalized area element $\Sigma_{\mu\nu}$. [^1] This is generic variation of our functional in the space of all functions $X^{A}_{\mu}(u,v)$, evaluated on our holomorphic solution after taking a variation. Had we computed the variation on the class of holomorphic functions by reducing the Dirichlet functional to the boundary, we would have found a symmetric quadratic form with a vanishing (and trivial) area derivative. Thus, variations orthogonal to the solution manifold are essential for the non-trivial area derivative. With our holomorphic solution, using the constant $B_{\mu\nu}$ tensor and the Dirichlet functional with the Hilbert transform, we find
@@
$\left.\frac{\delta^{2}S}{\delta C^{\prime}_{\mu}(\theta^{\prime})\delta C^{\prime}_{\nu}(\theta)}\right|_{\theta-0}^{\theta+0}\propto\int_{0}^{\theta}d\theta^{\prime}H(\theta^{\prime}-\theta)\hat{Q}_{\mu\nu};$ (6.6)
where $\hat{Q}_{\mu\nu}$ is the constant, antisymmetric tensor $\hat{Q}=\text{diag}(i\sigma_{2},i\sigma_{2})$. In the second derivative, by the chain rule, we use the standard Volterra derivative [Gelfand:1963:CalculusVariations] of the loop with respect to its tangent, $\delta C(x)/\delta C^{\prime}(y)=\Theta(x-y)$. This tensor $\hat{Q}$ is manifestly self-dual (as $\ast\hat{Q}_{12}=\hat{Q}_{34}$ and $\ast\hat{Q}_{34}=\hat{Q}_{12}$), which proves that the area derivative for $S_{\chi}[C;P]$ is self-dual. By the covariance argument above, this property extends to $S_{\mathrm{phys}}$.
Note that the numerical factor $\int_{0}^{\theta}d\theta^{\prime}H(\theta^{\prime}-\theta)\propto\sum_{n>0}1/n$ logarithmically diverges. However, the SD of the area derivative, valid for the tensor factor $Q_{\mu\nu}$, leads to the cancelation of the whole series in the loop space diffusion equation (2.5) (as argued in Sec. 2). Note that this cancelation is valid irrespective of the cutoff at large $n$ in Fourier series, so that if we fix this cutoff and substitute the area derivative into the loop equation (2.5), we get exactly zero. Thus, we can still claim the finite solution of the loop equation , as the minimal area is finite, and its contribution to the loop space diffusion equation is zero, making it a fixed point. ∎
@@ 7 The Hodge parity flip
How does our solution change for the negative Hodge parity $\chi=-1$ (ASD instead of SD)? There are many equivalent forms of the ASD solutions, all equivalent after the isometry transformation $O(4)$ rotation of the matrix $Q\Rightarrow QR$ in the Dirichlet BC $X=QC$. The conical transformation is a reflection $X^{A}_{4}\Rightarrow-X^{A}_{4}$ which in complex basis corresponds to $z_{2}\leftrightarrow\bar{z}_{2}$ in our holomorphic ansatz (4.3). After reflection, the only nonzero components will be
@@
As a result, the $(2,0)$ and $(0,2)$ components of the form $\Sigma$ will still be zero, and the $(0,1),(1,0)$ components change. Now the $\Sigma$ form will become $\Sigma=a(z)\tilde{B}$, where $\tilde{B}~e_{1}\wedge e_{2}-e_{4}\wedge e_{4}$ which has negative Hodge duality. The energy density $a(z)$ stays the same, and so does the area functional. The BC (4.5) change in an obvious way
@@
As expected, the area does not depend on the Hodge parity, but the area derivative switches from SD to ASD under Hodge parity flip. The loop equation will still hold, the same way as for the SD solution.
8 Reparametrization and additivity
Parametric invariance. As stated below (5.1) and in Sec. 3, the geometric area is defined by maximization over allowed isometries; we additionally take the infimum over orientation–preserving reparametrizations to remove any dependence on the boundary parameter. We therefore adopt the parametrization–invariant form
@@ 8 Reparametrization and additivity
Canonical split at a self–intersection. Let $C(\theta_{1})=C(\theta_{2})$ be a transversal self–intersection and define the two oriented subloops $C_{1}=C[\theta_{1},\theta_{2}]$ and $C_{2}=C[\theta_{2},\theta_{1}\!+\!2\pi]$ as in Sec. 3 (Case 2). Using parametric invariance, choose $\phi_{1},\phi_{2}\in\mathrm{Diff}^{+}(S^{1})$ with $\phi_{1}(S^{1})=[\theta_{1},\theta_{2}]$ and $\phi_{2}(S^{1})=[\theta_{2},\theta_{1}\!+\!2\pi]$ and set $\widetilde{C}_{j}(\tau)=C(\phi_{j}(\tau))$ so that each boundary component is on $0,2\pi]$. For a disconnected boundary the Douglas–Gram series splits,
@@
and, as explained in Sec. 3 (Case 2), the E–L problems decouple and the isometry group enlarges to the product $O(4)_{1}\times O(4)_{2}$.
Additivity of $S_{\mathrm{phys}}$. With (8.2), the double extremization factorizes:
@@
The first equality is the definition consistent with Sec. 3 (Case 2); the second uses the product structure of the objective (sum of two independent functionals), which makes the $\max$ over $R_{1},R_{2}$ and the $\min$ over $\phi_{1},\phi_{2}$ separate.
Compatibility with the MM kernel. The right–hand side of the MM equation [MMEq79, MM1981NPB] is supported on self–intersections via the bilocal kernel $\int\!\!\int d\theta_{1}d\theta_{2}\,C^{\prime}(\theta_{1})\!\cdot\!C^{\prime}(\theta_{2})\,\delta(C(\theta_{1})-C(\theta_{2}))\,\cdots$. Under $\theta_{a}=\phi_{a}(\tau_{a})$ (with $\phi_{a}^{\prime}\!>\!0$),
@@
So, the kernel is parameterization-invariant. Therefore, the canonical standardization to $0,2\pi]$ used above is legitimate and (8.4) guaranties that the multiplicative dressing $W\mapsto W\,\exp\left(-\kappa S\right)$ preserves the MM factorization $\mathcal{L}W[C_{1}\!\uplus\!C_{2}]=W[C_{1}]\otimes W[C_{2}]$ whenever $\mathcal{L}S=0$ (see (2.5) and [Mig83, migdal2025SQYMflow]).
@@ 9 Confinement theorem
Let $G$ be the Douglas–Gram matrix of $C$ with eigenvalues $0\leq\lambda_{1}\leq\cdots\leq\lambda_{4}$. as we have proven in the previous section, the Hodge dual area and the Euclidean minimal area satisfy
$A_{\mathrm{geom}}[C]:=\operatorname{tr}G\;=\;\sum_{i=1}^{4}\lambda_{i},\qquad S_{\chi}[C]\;=\;2\sqrt{2}\,(\lambda_{3}+\lambda_{4}),$ (9.1)
$\sigma_{\mathrm{eff}}[C]:=\frac{S_{\chi}[C]}{A_{\mathrm{geom}}[C]}\;=\;2\sqrt{2}\,\frac{\lambda_{3}+\lambda_{4}}{\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}}.$ (9.2)
@@
The quantity $\sigma_{\mathrm{eff}}[C]$ (hence $S_{\chi}[C]$) is O(4)O(4)–invariant and chirality–even: if $C\mapsto RC$ with $R\in O(4)$, then $G\mapsto RGR^{\top}$ leaves $\{\lambda_{i}\}$, and thus the theorem, unchanged.
@@
Both functionals, Douglas’s $\sum\lambda_{i}$ and our $\lambda_{3}+\lambda_{4}$, have yet to be minimized over the reparametrization of the bounding loop $C(\theta)\Rightarrow C(f(\theta)),f^{\prime}>0$. As the minimization of just the two highest eigenvalues, compared to that of the trace, is attained at different parameterizations, we can only state that $\sigma_{eff}=2\sqrt{2}\frac{\min(\lambda_{3}+\lambda_{4})}{\min(\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4})}$. However, the lower inequality still holds $\sigma_{eff}\geq\sqrt{2}$, as for every parametrization $\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}\leq 2(\lambda_{3}+\lambda_{4})$.Therefore, the minimum in the denominator is less than or equal to twice that in the numerator, proving the lower bound quoted in confinement theorem.
10 Master Field and the Meaning of the Minimal Surface in Planar QCD
Planar QCD represents the large-$N_{c}$ limit of a Quantum Field Theory defined by gauge fields, summing over all planar graphs to describe fluctuations around a (yet unknown) vacuum state. However, when formulated in terms of the Wilson loop observable $\Psi[C]=\langle\frac{1}{N}\mathrm{Tr}\mathcal{P}\exp\left(\oint_{C}Adx\right)\rangle$, the theory undergoes a fundamental shift: it ceases to be a quantum theory and becomes a classical field theory in loop space.
In the large $N_{c}$ limit, due to factorization $\langle W(C_{1})W(C_{2})\rangle=\langle W(C_{1})\rangle\langle W(C_{2})\rangle+O(1/N^{2})$, the functional $\Psi[C]$ becomes the Master Field, satisfying a deterministic nonlinear equation— the MM loop equation.
It is crucial to understand that in this framework, we are not “quantizing a string theory” in the traditional sense. Rather, we are solving a classical field equation in loop space. The introduction of the Hodge-dual minimal surface in the auxiliary space $\mathbb{R}^{4}\otimes\mathbb{R}^{4}$ is a method to construct this solution.
This approach mirrors the logic of the Eguchi-Kawai (EK) reduction [Das:1987dg]. In the EK model, quenched “momenta” or twisted boundary conditions are not introduced as inherent physical degrees of freedom of the vacuum, but as mathematical instruments designed to satisfy the loop equations of the planar theory on a single lattice site. Similarly, in our construction, the only physical requirement is the satisfaction of the MM equation; the string analogy applies strictly insofar as it serves this goal.
There is no claim of a direct physical connection between the microscopic QCD gluon variables and the embedding coordinates $X_{\mu}^{A}$ in the 16-dimensional auxiliary space. Unlike the AdS/CFT correspondence, we do not postulate a full quantum equivalence between planar QCD and a fundamental string theory in a higher-dimensional bulk. Instead, we construct the specific additive zero mode of the loop equation, in the same spirit as one solves for the spontaneous magnetization in a ferromagnet: identifying the non-perturbative classical configuration that creates the confining background.
Still, the technology of String Theory, if not its complete physical contents, can be quite useful in elaborating this solution of planar QCD.

### v3

@@ 6 Self-Duality of the Area Derivative
$\ast\frac{\delta S_{\mathrm{phys}}}{\delta\sigma_{\mu\nu}}=\chi\frac{\delta S_{\mathrm{phys}}}{\delta\sigma_{\mu\nu}}.$ (6.1)
@@
The physical action $S_{\mathrm{phys}}[\mathcal{C}]$ is the envelope (the maximum) of a family of functionals $S_{R}[C]:=S_{\chi}[RC;P]$, indexed by the rotation $R\in O(4)$, where $S_{\chi}[C;P]$ is the base functional from Sec. 4. The area derivative operator $\mathcal{D}:=\frac{\delta}{\delta\sigma}$ and the Hodge star $\ast$ are both $O(4)$-covariant. This covariance implies that if we can prove self-duality for any single member of this family (e.g., the $R=I$ case, $S_{\chi}[C;P]$), then self-duality holds for all members. The derivative of the envelope $S_{\mathrm{phys}}$ at a loop $C$ is, by definition, the derivative of the specific functional $S_{R_{\text{max}}}[C]$ that achieves the maximum at that point. Since all functionals in the family will be shown to have an SD derivative, $S_{\mathrm{phys}}$ must also have an SD derivative. Therefore, the proof reduces to verifying the self-duality for the base functional $S_{\chi}[C;P]$ (for $R=I$), which we do now.
@@
where $T_{\mu\nu}$ is the normalized area element $\Sigma_{\mu\nu}$. With our holomorphic solution, using the constant $B_{\mu\nu}$ tensor and the Dirichlet functional with the Hilbert transform, we find
@@
$\frac{\delta^{2}S}{\delta C^{\prime}_{[\mu}(\theta-0)\delta C^{\prime}_{\nu]}(\theta)}\propto\oint d\theta^{\prime}H(\theta^{\prime}-\theta)\Theta(\theta^{\prime}-\theta)\hat{Q}_{\mu\nu};$ (6.6)
where $\hat{Q}_{\mu\nu}$ is the ’t Hooft SD or ASD tensor $\hat{Q}_{\mu\nu}=\eta^{3}_{\mu\nu}\text{ or }\bar{\eta}^{3}_{\mu\nu}$. In the second derivative, by the chain rule, we use the standard Volterra derivative [Gelfand:1963:CalculusVariations] of the loop with respect to its tangent, $\delta C(x)/\delta C^{\prime}(y)=\Theta(x-y)$. This tensor $\hat{Q}$ is manifestly self-dual (as $\ast\hat{Q}_{12}=\chi\hat{Q}_{34}$ and $\ast\hat{Q}_{34}=\chi\hat{Q}_{12}$), which proves that the area derivative for $S_{\chi}[C;P]$ is (anti)self-dual. By the covariance argument above, this property extends to $S_{\mathrm{phys}}$.
Note that the numerical factor $\oint d\theta^{\prime}H(\theta^{\prime}-\theta)\Theta(\theta^{\prime}-\theta)\propto\sum_{\text{odd }n>0}1/n$ logarithmically diverges. However, the SD of the area derivative, valid for the tensor factor $Q_{\mu\nu}$, leads to the cancellation of the whole series in the loop space diffusion equation (2.5) (as argued in Sec. 2). Note that this cancellation is valid irrespective of the cutoff at large $n$ in the Fourier series, so that if we fix this cutoff and substitute the area derivative into the loop equation (2.5), we get exactly zero. Thus, we can still claim a finite solution to the loop equation, since the minimal area is finite and its contribution to the loop space diffusion equation is zero, making it a fixed point. ∎
Remark 6.1.
In the computation of the area derivative we use a generic variation of the functional in the full space of embeddings $X^{A}_{\mu}(u,v)$, evaluated on the holomorphic solution after taking the variation. Had we restricted the variation to the holomorphic submanifold $X^{A}_{\mu}(u,v)=\Re\,U^{A}_{\mu}(z)$ (preserving the complex structure), we would have found a vanishing area derivative, since the extremal area is stationary with respect to variations of the moduli. Thus, variations orthogonal to the solution manifold are essential for obtaining a nontrivial area derivative. It is precisely through these orthogonal variations that the boundary condition X=QCX=Q\,C (which carries the Hodge-parity label $\chi$) enters the variation: for $\chi=+1$ the resulting area derivative is self-dual, whereas for $\chi=-1$ it is anti-self-dual, even though the on-shell scalar value $S_{\chi}[C]=2\sqrt{2}\,(\lambda_{3}+\lambda_{4})$ is the same in both cases.
A simple analogy is a family of functions $F(x,y;\chi)$ that share a degenerate extremal curve $(x_{0}(t),y_{0}(t))$, such that the extremal value $F(x_{0}(t),y_{0}(t);\chi)$ is independent of $\chi$ for any $t$, while the Hessian components orthogonal to this curve, $(\partial_{\perp})^{2}F$, evaluated on $(x_{0}(t),y_{0}(t))$ still depend on $\chi$. Likewise, $S_{\chi}[C]$ has the same extremal value for both branches, but its area derivative in loop space retains the SD/ASD label. There is therefore no contradiction between the $\chi$-independence of the minimal area and the $\chi$-dependence of its area derivative, which is the antisymmetric two-form part of the second functional derivative of $S[C]$ with respect to the loop tangent $C^{\prime}(\theta)$.
@@ 7 The Hodge parity flip
How does our solution change for the negative Hodge parity $\chi=-1$ (ASD instead of SD)? There are many equivalent forms of the ASD solutions, all equivalent after the isometry transformation $O(4)$ rotation of the matrix $Q\Rightarrow QR$ in the Dirichlet BC $X=QC$. The canonical transformation is a T-parity: reflection $X^{A}_{4}\Rightarrow-X^{A}_{4}$ which in complex basis corresponds to $z_{2}\leftrightarrow\bar{z}_{2}$ in our holomorphic ansatz (4.3). After reflection, the only nonzero components will be
@@
As a result, the $(2,0)$ and $(0,2)$ components of the form $\Sigma$ will still be zero, and the $(0,1),(1,0)$ components change. Now the $\Sigma$ form will become $\Sigma=a(z)\tilde{B}$, where $\tilde{B}=e_{1}\wedge e_{2}-e_{4}\wedge e_{4}$ which has negative Hodge duality. The energy density $a(z)$ remains the same, and so does the area functional. The BC (4.5) change in an obvious way
@@
As expected, the scalar value of the area does not depend on the Hodge parity $\chi$, as it is determined solely by the parity-invariant spectrum of the Douglas-Gram matrix. However, the area derivative—the ‘force’ term entering the loop equation—transforms non-trivially, switching from Self-Dual (SD) to Anti-Self-Dual (ASD) under the parity flip. Since the loop diffusion operator $\mathcal{L}$ annihilates both SD and ASD tensor structures independently via the Bianchi identity, both branches $\chi=\pm 1$ represent valid, degenerate zero modes of the equation. To construct a fully Parity-invariant solution for the Wilson loop in QCD, we must treat these degenerate sectors symmetrically. The physical dressing factor involves the sum of both branches in the exponent:
$\exp\left(-\frac{\kappa}{2}\left(S_{\chi=+1}[C]+S_{\chi=-1}[C]\right)\right).$
Since the extremal area values are identical, $S_{+}=S_{-}=S$, this sum restores the discrete symmetry of the theory without altering the spectral form of the confining area law.
8 Reparametrization and factorization for the mulri-loop hierarchy
8.1 Parametric invariance.
As stated below (5.1) and in Sec. 3, the geometric area is defined by maximization over allowed isometries; we additionally take the infimum over orientation–preserving reparametrizations to remove any dependence on the boundary parameter. We therefore adopt the parametrization–invariant form
@@ 8 Reparametrization and factorization for the mulri-loop hierarchy
8.2 Canonical split at a self-intersection.
Let $C(\theta_{1})=C(\theta_{2})$ be a transversal self–intersection and define the two oriented subloops $C_{1}=C[\theta_{1},\theta_{2}]$ and $C_{2}=C[\theta_{2},\theta_{1}\!+\!2\pi]$ as in Sec. 3 (Case 2). Using parametric invariance, choose $\phi_{1},\phi_{2}\in\mathrm{Diff}^{+}(S^{1})$ with $\phi_{1}(S^{1})=[\theta_{1},\theta_{2}]$ and $\phi_{2}(S^{1})=[\theta_{2},\theta_{1}\!+\!2\pi]$ and set $\widetilde{C}_{j}(\tau)=C(\phi_{j}(\tau))$ so that each boundary component is on $0,2\pi]$. For a disconnected boundary, the Douglas–Gram series splits,
@@
and, as explained in Sec. 3 (Case 2), the E–L problems decouple, and the isometry group enlarges to the product $O(4)_{1}\times O(4)_{2}$.
8.3 Additivity of $S_{\mathrm{phys}}$.
With (8.2), the double extremization factorizes:
@@
The first equality is the definition consistent with Sec. 3 (Case 2); The second one uses the product structure of the objective (sum of two independent functionals), which makes the $\max$ over $R_{1},R_{2}$ and the $\min$ over $\phi_{1},\phi_{2}$ separate.
8.4 Compatibility with the MM kernel.
The right–hand side of the MM equation [MMEq79, MM1981NPB] is supported on self–intersections via the bilocal kernel $\int\!\!\int d\theta_{1}d\theta_{2}\,C^{\prime}(\theta_{1})\!\cdot\!C^{\prime}(\theta_{2})\,\delta(C(\theta_{1})-C(\theta_{2}))\,\cdots$. Under $\theta_{a}=\phi_{a}(\tau_{a})$ (with $\phi_{a}^{\prime}\!>\!0$),
@@
So, the kernel is parameterization-invariant. Therefore, the canonical standardization to $0,2\pi]$ used above is legitimate and (8.4) guaranties that the multiplicative dressing $W\mapsto W\,\exp\left(-\kappa S\right)$ preserves the MM factorization $\mathcal{L}W[C_{1}\!\circ\!C_{2}]=W[C_{1}]\otimes W[C_{2}]$ whenever $\mathcal{L}S=0$ (see (2.5) and [Mig83, migdal2025SQYMflow]).
@@
8.5 Confining factor in the multiloop hierarchy
Lemma 8.1 (Finite-$N_{c}$ dressing symmetry).
Let $\{W_{n}\}_{n\geq 1}$ denote the $n$-loop Wilson loop averages in pure $\mathrm{SU}(N_{c})$ Yang–Mills theory, satisfying the MM multiloop hierarchy [Mig83]. Let $S[C]$ be the geometric functional constructed above, with the properties
  1. 1.
    $LS=0$, where $L$ is the loop diffusion operator (2.5);
  2. 2.
    strict additivity at intersections: whenever a loop $C$ self-intersects at $x=y$ into subloops $C_{xy}$ and $C_{yx}$, one has
    $S[C]=S[C_{xy}]+S[C_{yx}],$
    and similarly for intersections between distinct loops.
For any real parameter $\kappa$, define the multiplicatively dressed correlators
$\widetilde{W}_{n}(C_{1},\dots,C_{n})$
$:=W_{n}(C_{1},\dots,C_{n})\,\exp\left(-\kappa\sum_{i=1}^{n}S[C_{i}]\right).$ (8.6)
Then, for any finite $N_{c}$, the family $\{\widetilde{W}_{n}\}$ also satisfies the full MM multiloop hierarchy with the same $N_{c}$.
Proof.
The same loop operator $L$ in (2.5) appears on the left-hand side of the general QCD loop equations [Mig83] for arbitrary $N_{c}$. The difference between the planar and finite-$N_{c}$ cases lies only in the right-hand side: at a self-intersection, the planar factorization is replaced by a combination of one- and two-loop averages, but the kernel of the bilocal integral is the same.
At a self-intersection point of a single loop $C$, the finite-$N_{c}$ loop equation involves the combination:
$W_{2}[C_{xy},C_{yx}]-\frac{1}{N_{c}^{2}}\,W_{1}[C_{xy}\circ C_{yx}].$
This equation is one link in the chain relating $W_{n}$ to $W_{n\pm 1}$; for finite $N_{c}$ the chain terminates due to the Fierz identity for $\mathrm{SU}(N_{c})$ [Mig83], a fact recently exploited in lattice studies of the loop equation [kazakovZheng2024].
In the general multiloop case, $W_{n}[C^{1},\dots,C^{n}]$, the hierarchy contains integrals over pairs of points $x,y$ lying on two (not necessarily distinct) loops $C^{i},C^{j}$, again with the standard MM kernel. When $x$ and $y$ lie on two distinct loops $C^{i}$ and $C^{j}$, the corresponding term involves a merging of these loops in $W_{n-1}$ at the intersection point,
$\frac{1}{N_{c}^{2}}\,W_{n-1}[\dots,C^{i}_{xy}\circ C^{j}_{yx},\dots]\;-\;W_{n}[\dots,C^{i}_{xy},C^{j}_{yx},\dots].$
Thus, every equation in the hierarchy relates loop functionals evaluated on collections of loops that differ only by local splittings or mergings at intersection points.
Now substitute the dressing ansatz (8.6) into the multiloop hierarchy. Because $\mathcal{L}S=0$ and $\mathcal{L}$ is linear, the loop operator on the left-hand side passes through the exponential factor and acts only on the perturbative part:
$\mathcal{L}\,\widetilde{W}_{n}(C_{1},\dots,C_{n})$
$=\exp\left(-\kappa\sum_{i}S[C_{i}]\right)\,\mathcal{L}\,W_{n,\mathrm{pert}}(C_{1},\dots,C_{n}).$ (8.7)
In particular, the dressed correlators satisfy the same differential operator $\mathcal{L}$ as the undressed ones.
On the right-hand side, consider first a self-intersection of a single loop $C$. After dressing, the corresponding contribution becomes
$\Big(W_{2,\mathrm{pert}}(C_{xy},C_{yx})-\frac{1}{N_{c}^{2}}W_{1,\mathrm{pert}}(C_{xy}\circ C_{yx})\Big)$
$\exp\left(-\kappa(S[C_{xy}]+S[C_{yx}])\right).$ (8.8)
By strict additivity, $S[C]=S[C_{xy}]+S[C_{yx}]$ at the intersection point $x=y$. Thus, the exponential factor in (8.8) is the same as the factor multiplying the left-hand side, and therefore cancels from the equation.
For an intersection of two distinct loops $C^{i}$ and $C^{j}$, the right-hand side contains terms of the schematic form
$\Big(W_{n,\mathrm{pert}}(C^{i}_{xy},C^{j}_{yx})-\frac{1}{N_{c}^{2}}\,W_{n-1,\mathrm{pert}}(C^{i}_{xy}\circ C^{j}_{yx})\Big)$
$\times\exp\left(-\kappa\big(S[C^{i}_{xy}]+S[C^{j}_{yx}]+\sum_{k\neq i,j}S[C^{k}]\big)\right).$ (8.9)
Again, by additivity at the intersection and for disjoint unions,
$S[C^{i}]+S[C^{j}]$
$=S[C^{i}_{xy}]+S[C^{i}_{yx}]+S[C^{j}_{xy}]+S[C^{j}_{yx}]$
$=S[C^{i}_{xy}\uplus C^{j}_{yx}]+S[C^{j}_{xy}\uplus C^{i}_{yx}]$
$=S[C^{i}_{xy}\uplus C^{j}_{yx}\uplus C^{j}_{xy}\uplus C^{i}_{yx}]$ (8.10)
so the sum $\sum_{i}S[C^{i}]$ is invariant under the local splitting/merging operation. Hence, the same exponential factor multiplies every term in the multiloop equation and cancels between the two sides.
Since this reasoning applies to each equation in the finite-$N_{c}$ hierarchy and uses only (i) the universal kernel and operator $\mathcal{L}$ and (ii) the additivity and zero-mode properties of $S[C]$, it follows that the dressed correlators $\widetilde{W}_{n}$ satisfy the same multiloop MM hierarchy for any finite $N_{c}$. ∎
Remark 8.1.
Note that the perimeter of the set of loops is also an additive functional, factorizing over the right side of the MM chain of equations. However, this is not a Stokes-type functional, so it does not process finite area derivatives, and, as a consequence, it cannot be added to the zero mode of the loop operator $\mathcal{L}$. Such terms arise as quark-mass renormalization corrections in the perturbative Wilson-loop factor; they are linearly divergent and depend on the regularization scheme. In any case, these terms do not affect the area law at large loops, which is responsible for quark confinement.
@@ 9 Confinement theorem
Let $G$ be the Douglas–Gram matrix of a regular loop $C$ with eigenvalues $0\leq\lambda_{1}\leq\cdots\leq\lambda_{4}$. As proven in the previous sections, the Euclidean geometric area and the Hodge–dual area satisfy
$A_{\mathrm{geom}}[C]:=\operatorname{tr}G\;=\;\sum_{i=1}^{4}\lambda_{i},\hskip 18.49988ptS[C]\;=\;2\sqrt{2}\,(\lambda_{3}+\lambda_{4}),$ (9.1)
$\sigma_{\mathrm{eff}}[C]:=\frac{S[C]}{A_{\mathrm{geom}}[C]}\;=\;2\sqrt{2}\,\frac{\lambda_{3}+\lambda_{4}}{\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}}.$ (9.2)
@@
The quantity $\sigma_{\mathrm{eff}}[C]$ (hence $S[C]$) is O(4)O(4)–invariant and chirality–even: if $C\mapsto RC$ with $R\in O(4)$, then $G\mapsto RGR^{\top}$ leaves $\{\lambda_{i}\}$, and thus the theorem, unchanged. Combined with Lemma 8.1, this implies that the multiplicative factor $\exp\left(-\kappa S[C]\right)$ defines an exact confining area law solution of the MM loop equations for pure $\mathrm{SU}(N_{c})$ Yang–Mills theory with an arbitrary number of colors $N_{c}$, not just in the planar limit $N_{c}=\infty$.
@@
Both functionals, Douglas’s $\sum_{i}\lambda_{i}$ and our $\lambda_{3}+\lambda_{4}$, must still be minimized over the reparametrization of the bounding loop, $C(\theta)\mapsto C(f(\theta))$ with $f^{\prime}>0$. Since the minimization of just the two highest eigenvalues, as opposed to that of the trace, is generally attained at a different parametrization, we can only state that
$\sigma_{\mathrm{eff}}[C]=2\sqrt{2}\,\frac{\min(\lambda_{3}+\lambda_{4})}{\min(\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4})}.$
However, the lower inequality still holds: $\sigma_{\mathrm{eff}}[C]\geq\sqrt{2}$, because for every parametrization one has $\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}\leq 2(\lambda_{3}+\lambda_{4})$. Therefore, the minimum in the denominator is less than or equal to twice the minimum in the numerator, which proves the lower bound quoted in the Confinement Theorem.